# Evaluating Evaluation Metrics – The Mirage of Hallucination Detection

### v1

@@ 1 Introduction
Hallucinations in language model generations are detrimental and, unfortunately, a pervasive phenomenon Kaddour et al. (2023); Xu et al. (2024). They manifest as texts that are nonsensical, degenerate, factually incorrect, or unverifiable based on the input context Ji et al. (2023). As language models are rapidly adopted across various settings, addressing hallucinations has become a key research focus Varshney et al. (2023); Dhuliawala et al. (2023); Chuang et al. (2024); Shi et al. (2024). However, before investing time and resources into devising its mitigation techniques, it is worthwhile to take a step back and ask: 1) Are the existing metrics truly capturing the hallucinations effectively? 2) Do these metrics generalize across different datasets, decoding techniques, model families, and model sizes? Confronting these questions is vital, as any attempt to alleviate hallucinations is futile unless we ensure its robust, reliable, and accurate measurement.
Moreover, the term ‘hallucination’ covers a spectrum of generation errors. In this work, we focus on its two most common manifestations: poor faithfulness and factuality, particularly in knowledge-grounded dialog Dziri et al. (2022a, b) and question-answering Lin et al. (2022); Li et al. (2023a). Faithfulness measures the consistency and truthfulness with the provided knowledge source, while factuality pertains to the accuracy with respect to real-world facts or widely accepted knowledge Maynez et al. (2020).
@@ 2 Experimental Setup
More details on datasets, models, and metrics are provided in Appendix §A.1, A.2, and A.3, respectively.
@@
We focus on four datasets. FaithDial Dziri et al. (2022a) and Begin Dziri et al. (2022b) are knowledge-grounded dialog datasets, where, given a conversation history $H=(u_{1},\dots,u_{n-1})$ and knowledge source $K_{n}$, the system must generate a response $\bar{u}_{n}$ that is coherent with $H$ and supported by a non-empty subset $M_{n}\subset K_{n}$ to be considered faithful. TruthfulQA Lin et al. (2022) is a factual question-answering dataset with multiple plausible answers. We measure factuality by comparing the generated answer’s alignment with them. Lastly, we analyze the knowledge-grounded QA and dialog subsets of the HaluEval Li et al. (2023a) benchmark. Begin and HaluEval already contain responses from pretrained models and ChatGPT, respectively, annotated for hallucinations. We generate responses for TruthfulQA and FaithDial.
@@
Our study includes five language model families: OPT Zhang et al. (2022), Llama Touvron et al. (2023); Dubey et al. (2024), OLMo Groeneveld et al. (2024), Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024), and Gemma Gemma (2024). We cover models ranging from 125M to 70B, including their instruction-tuned versions, totaling 37 models. Evaluation spans five decoding methods of greedy, beam search Graves (2012), ancestral, top-k Fan et al. (2018), and top-p sampling Holtzman et al. (2020).
@@
Begin CMU 0.77 ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.84}}}$ $0.65$ $0.70$ $0.73$ $0.65$
Begin TC 0.74 $0.71$ $0.65$ $0.67$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.76}}}$ $0.65$
Begin WoW 0.83 $0.77$ $0.43$ $0.43$ $0.56$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.96}}}$
HaluEval Dial $0.49$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.74}}}$ $0.39$ 0.61 $0.54$ $0.42$
HaluEval QA $0.53$ $0.66$ $0.36$ 0.83 $0.82$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.93}}}$
Average $0.67$ ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.74}}}$ $0.50$ $0.65$ $0.68$ 0.72
@@
3 Results and Discussion
Table 1 displays the alignment scores of various metrics with human labels. Using PRAUC for continuous metrics and weighted-F1 for binary metrics (with random baseline scores of $0.50$ and $0.50-0.56$, respectively), we find mixed results across evaluation methods. The UniEval suite’s factual consistency evaluator performed just about at or below random chance across all the six data subsets. K-BertScore and $\mathcal{Q}^{2}$ NLI both show strong performance on Begin CMU and HaluEval QA, with the latter also doing well on Begin TC. However, they both struggle to replicate performance on Begin WoW and HaluEval Dial. Critic, as expected, excels on the Begin benchmark, since it is trained on dialog datasets. However, surprisingly, it drastically under performs on the HaluEval tasks, faring worse than even the random baseline. The GPT-4 evaluator consistently shows agreeable alignment on average, acing on two datasets. Our proposed ensemble metric is a close second, excelling particularly on the Begin WoW and HaluEval QA subsets. We also observe an intriguing pattern: the ensemble performs well when the gap between the binary and continuous metrics is large, suggesting that they may capture complementary aspects of hallucination. We discuss more in Appendix §C.1.
@@ 3 Results and Discussion
Finding 3 Instruction tuning and mode-seeking decoding methods reduce hallucinations
@@
Table 2: Significance test results (rounded to 3 decimals) for the impact of training type, model size, and decoding methods on hallucination metrics. Red cells ($p>0.05$) indicate failure to reject the null hypothesis.
Figure 3: Hallucination detection metric scores for greedy decoding as model size increases. Circles and hexagons represent pretrained and instruction-tuned models, respectively.
@@
Instruction-tuning is known to perform well on grounded generation tasks, and to reduce hallucinations Ouyang et al. (2022); Dhuliawala et al. (2023); Kalai and Vempala (2024). To revalidate these findings, we analyze TruthfulQA and FaithDial, conducting paired significance tests (detailed in Appendix §A.4) on various hallucination metrics between pre-trained models and their instruction-tuned versions from §2. The null hypothesis posits that instruction-tuning has no effect on these metrics. The results in Table 2 help us refute this claim, albeit with some exceptions, where SacreBleu and K-BertScore for FaithDial show no significant gains with instruction-tuning. Nevertheless, the null hypothesis is rejected for the more reliable metrics of Critic and GPT-4, suggesting that instruction-tuning effectively reduces hallucinations.
Shifting focus to decoding techniques, it is well established that mode-seeking decoding methods such and greedy and beam search tend to hallucinate less than sampling methods (ancestral, top-p, and top-k) Dziri et al. (2021); Li et al. (2024). Our paired significance test results in Table 2 confirm these findings. Additionally, the posthoc pairwise significance testing results in Figures 10 and 11 (Appendix §C.2) strengthen our argument.
@@
Finding 4 Parameter scaling doesn’t lead to commensurate gains across all metrics and datasets
@@
Scaling language model parameters typically leads to a monotonic increase in both pretraining Kaplan et al. (2020); Hoffmann et al. (2022) and downstream metrics Caballero et al. (2023), often following a power law. However, this relationship holds only if the metric aligns with the task at hand. Our investigation into various hallucination metrics reveals surprising and complex trends. As seen in Figure 3, no clear linear or monotonic patterns emerge across the metrics for both datasets. Critic also shows contradictory trends in TruthfulQA and Faithdial. Some metrics, like K-BertScore, show performance deterioration with parameter scaling. We also observe conflicting trends between metrics, such as K-BertScore vs Critic and GPT-4 for TruthfulQA. The results for Gemma 1 and 1.1 often suggest opposite conclusions regarding hallucinations. Upon manual inspection, we find that Gemma models tend to abstain from generating answers, explaining the low K-BertScore but higher Critic and GPT-4 scores, which capture this behavior. Similar underperformance trends are evident across other metric types, as shown in Figures 12 and 13. More findings are presented in Appendix §C.3.
@@
For further analysis, we bin models by size and perform an unpaired statistical test. The null hypothesis is that parameter scaling has no effect on metric performance. As shown in Table 2, only GPT-4 consistently rejects the null hypothesis, indicating its performance improves with size. Figure 4 shows posthoc pairwise p-values. $\mathcal{Q}^{2}$ NLI and Critic for FaithDial, and K-BertScore for TruthfulQA, show little improvement with parameter scaling. Overall, increasing model size does not always improve downstream hallucination detection metrics.
@@ 4 Conclusion
Hallucination detection in natural language generation is a tricky task. Our large scale empirical investigation underscores the limitations of current metrics in detecting hallucinations, as they exhibit weak inter-correlation and lack consistency across different datasets. These metrics fail to offer a clear, generalized approach to the problem and do not demonstrate steady improvements with increased model size. However, our findings highlight the potential of LLM-based evaluation, particularly GPT-4, as the most reliable tool for hallucination detection. Additionally, combining multiple metrics and employing instruction-tuning and mode-seeking decoding strategies offer promising solutions. Ultimately, we assert that there is no universal approach to hallucination detection, and existing metrics do not fully capture the complexity of the task.
Limitations
While our large-scale empirical investigation provides a thorough analysis of the current hallucination metrics, it does have certain limitations. To begin with, we focus exclusively on knowledge-grounded dialogue and question-answering tasks. However, hallucination is a prevalent issue across various other NLP tasks, such as machine translation Raunak et al. (2021), summarization, and linguistic applications Weissweiler et al. (2023). Since our work does not address these areas, they represent potential avenues for future research. Additionally, while fine-tuning has been shown to mitigate model hallucinations Ghosal et al. (2024), we have not explored these experiments in our study, leaving them for future investigation.
@@ A.1 Datasets
  • Begin Benchmark Dziri et al. (2022b) It is a collection of 3 knowledge-grounded dialog datasets – CMU-Dog Zhou et al. (2018), Wizard of Wikipedia (WoW) Dinan et al. (2019), and TopicalChat Gopalakrishnan et al. (2019). It contains responses generated by 4 models – GPT2 Radford et al. (2019), T5 Raffel et al. (2020), DoHA (BART with dual attention) Prabhumoye et al. (2021), and CTRL-T5 (control tokens augmented T5) Rashkin et al. (2021). Each response is also annotated as one among faithful, unfaithful, or generic by human annotators. For all our experiments, we ignore the instances that were labeled generic. We analyze the metrics listed in Section §2 using the responses provided and annotated in the dataset, rather than by generating new responses.
  • HaluEval Li et al. (2023a) It is a conglomerate of $5,000$ general-purpose and $30,000$ task-specific examples designed for hallucination evaluation, spanning question answering, knowledge-grounded dialog, and text summarization. We focus on the task-specific subset, which includes $10,000$ examples randomly sampled from the training sets of HotpotQA Yang et al. (2018), and OpenDialKG Moon et al. (2019). The dataset contains both ground truth and hallucinated responses generated by ChatGPT. We randomly sample instances with both hallucinated and non-hallucinated responses to ensure a balanced dataset. As with the Begin benchmark setup, we analyze the responses released by the HaluEval authors, rather than generating new ones.
  • TruthfulQA Lin et al. (2022) It assess how accurately a language model generates truthful answers to factual questions. It features $817$ questions across $38$ diverse categories. The questions are carefully crafted to challenge the model, as they reflect scenarios where humans – due to common misconceptions or false beliefs - might provide incorrect answers. It contains two setups – text generation and multiple choice questions. We consider the former in our experiments. We generate the answers using the models mentioned in §2.
  • FaithDial Dziri et al. (2022a) It is an information-seeking, knowledge-grounded dialog response generation dataset, that contains $50,000$ dialog turns spread across $5,500$ conversations. It is curated by refining and correcting the hallucinated responses found in the Wizard-of-Wikipedia (WoW) corpus Dinan et al. (2019). We generate the responses using the models mentioned in §2.

### v2

@@ 1 Introduction
Hallucinations in language model generations are detrimental and, unfortunately, a pervasive phenomenon Ji et al. (2023); Kaddour et al. (2023); Xu et al. (2024b). As language models are rapidly adopted across various settings, addressing hallucinations has become a key research focus Varshney et al. (2023); Dhuliawala et al. (2023); Chuang et al. (2024); Shi et al. (2024). However, before investing time and resources into devising its mitigation techniques, it is worthwhile to take a step back and ask: 1) Are the existing metrics truly capturing the hallucinations effectively? 2) Do these metrics generalize across different datasets, decoding techniques, model families, and model sizes? Confronting these questions is vital, as any attempt to alleviate hallucinations is futile unless we ensure its robust, reliable, and accurate measurement.
The term ‘hallucination’ covers a spectrum of generation errors. In this work, we focus on its two most common manifestations: poor faithfulness and factuality, particularly in knowledge-grounded dialog Dziri et al. (2022a, b) and question-answering Lin et al. (2022); Li et al. (2023a). Faithfulness measures the consistency and truthfulness with the provided knowledge source, while factuality pertains to the accuracy with respect to real-world facts or widely accepted knowledge Maynez et al. (2020).
@@ 2 Experimental Setup
We focus on four datasets. FaithDial Dziri et al. (2022a) and Begin Dziri et al. (2022b) are knowledge-grounded dialog datasets, where, given a conversation history $H=(u_{1},\dots,u_{n-1})$ and knowledge source $K_{n}$, the system generates a response $\bar{u}_{n}$ that is coherent with $H$ and supported by a non-empty subset $M_{n}\subset K_{n}$ to be considered faithful. TruthfulQA Lin et al. (2022) is a factual question-answering dataset with multiple plausible answers. We measure factuality by comparing the generated answer’s alignment with them. Lastly, we analyze the knowledge-grounded QA and dialog subsets of the HaluEval Li et al. (2023a) benchmark. More details are provided in Appendix §A.1.
@@
Our study includes five LLM families: OPT Zhang et al. (2022), Llama Touvron et al. (2023); Dubey et al. (2024), OLMo Groeneveld et al. (2024), Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024), and Gemma Gemma (2024). We cover models ranging from 125M to 70B, including their instruction-tuned versions, totaling 37 models. Evaluation spans five decoding methods of greedy, beam search Graves (2012), ancestral, top-k Fan et al. (2018), and top-p sampling Holtzman et al. (2020).
@@ 3 Results and Discussion
[…]
Begin TC 0.74 $0.71$ $0.65$ $0.76¯{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underline{\textbf{0.76}}}$ $0.65$
[…]
@@
Table 1 displays the alignment scores of various metrics with human labels.  [^1] Only Begin and HaluEval are included in this study, as they provide LLM-generated responses labeled for hallucinations.. Using PRAUC for continuous metrics and weighted-F1 for binary metrics (with random baseline scores of $0.50$ and 0.500.560.50-0.56, respectively), we find mixed results across evaluation methods. The UniEval suite’s factual consistency evaluator performed just about at or below random chance across all the six data subsets. K-BertScore and $\mathcal{Q}^{2}$ NLI both show strong performance on Begin CMU and HaluEval QA, with the latter also doing well on Begin TC. However, they both struggle to replicate performance on Begin WoW and HaluEval Dial. Critic, as expected, excels on the Begin benchmark, since it is trained on dialog datasets. However, surprisingly, it drastically under performs on the HaluEval tasks, faring worse than even the random baseline. The GPT-4 evaluator consistently shows agreeable alignment on average, acing on two datasets. Our proposed ensemble metric is a close second, excelling particularly on the Begin WoW and HaluEval QA subsets. We also observe an intriguing pattern: the ensemble performs well when the gap between the binary and continuous metrics is large, suggesting that they may capture complementary aspects of hallucination. We discuss more metrics and related findings in Appendix C.1.
@@
Finding 3 Instruction-tuning and mode-seeking decoding methods reduce hallucinations
@@
Table 2: Significance test results for the impact of training type, model size, and decoding methods on hallucination metrics. Red cells ($p>0.05$) indicate failure to reject the null hypothesis.
Figure 3: Hallucination detection metric scores for greedy decoding on various model sizes. Circles and hexagons represent pretrained and instruction-tuned models, respectively.
@@
Instruction-tuning is known to perform well on grounded generation tasks, and to reduce hallucinations Ouyang et al. (2022); Dhuliawala et al. (2023); Kalai and Vempala (2024). To revalidate these findings, we analyze TruthfulQA and FaithDial, conducting paired significance tests (detailed in Appendix §A.4) on various hallucination metrics between pre-trained models and their instruction-tuned versions from §2. The null hypothesis posits that ‘Instruction-tuning has no effect on hallucination detection metrics’. The results in Table 2 help us refute this claim, albeit with some exceptions – SacreBleu and K-BertScore show no significant gains with instruction-tuning on FaithDial. Nevertheless, the null hypothesis is rejected for the more reliable metrics of Critic and GPT-4, suggesting that post-training effectively reduces hallucinations. Shifting focus to decoding techniques, it is well established that mode-seeking decoding methods such and greedy and beam search tend to hallucinate less than sampling methods (ancestral, top-p, and top-k) Dziri et al. (2021); Li et al. (2024). Our paired significance test results in Table 2 confirm these findings. Additionally, the posthoc pairwise significance testing results in Figures 11 and 12 (Appendix §C.3) strengthen our argument.
@@
Finding 4 Metrics do not show commensurate gains with parameter scaling
@@
Scaling language model parameters typically leads to a monotonic increase in both pretraining Kaplan et al. (2020); Hoffmann et al. (2022) and downstream metrics Caballero et al. (2023), often following a power law. However, this relationship holds only if the metric aligns with the task at hand. Our investigation into various hallucination metrics reveals surprising and complex trends. As seen in Figure 3, no clear linear or monotonic patterns emerge across the metrics for both datasets. Critic also shows contradictory trends in TruthfulQA and Faithdial. Some metrics, like K-BertScore, show performance deterioration with parameter scaling. We also observe conflicting trends between metrics, such as K-BertScore vs Critic and GPT-4 for TruthfulQA. The results for Gemma 1 and 1.1 often suggest opposite conclusions regarding hallucinations. Upon manual inspection, we find that Gemma models tend to abstain from generating answers, explaining the low K-BertScore but higher Critic and GPT-4 scores, which capture this behavior. Similar underperformance trends are evident across other metric types, as shown in Figures 13 and 14. More findings are presented in Appendix §C.4.
@@
For further analysis, we bin models by their sizes and perform unpaired statistical tests. The null hypothesis here is that ‘Parameter scaling has no effect on metric performance’. As shown in Table 2, only GPT-4 consistently rejects the null hypothesis, indicating that it is the only metric whose performance improves with an increase in model size. Figure 4 shows posthoc pairwise p-values. $\mathcal{Q}^{2}$ NLI and Critic for FaithDial, and K-BertScore for TruthfulQA, show little improvement with parameter scaling. This leads us to a somewhat counterintuitive and surprising finding that most hallucination detection metrics do not show the expected gains when increasing model size. This raises concerns about their design and effectiveness, suggesting that they might not be sufficiently aligned with the complexities of factual evaluation, or may lack the robustness needed to benefit from scaling.
4 Failure Modes of the Metrics
Figure 5: Metric accuracy across varying response lengths on the Begin and HaluEval.
To better understand the limitations and bottlenecks of evaluation metrics, we analyze how their performance varies with respect to instance difficulty. While ‘difficulty’ can be quantified in various ways, it is well-established that LLMs tend to struggle and exhibit increased hallucination rates, with longer responses Yang et al. (2025); ul Islam et al. (2025). Thus, we investigate the potency of different metrics across varying response lengths. Specifically, we calculate the accuracy of each metric for different response lengths, as shown in Figure 5. Appendix C contains additional error analysis and discussion.
UniEval suite’s Consistency performs reasonably well on longer responses in the Begin CMU and Begin TC datasets, where accuracy rises steadily with response length – reaching values above $0.9$. However, its performance on Begin WOW starts quite low ($0.4$ in short completions), and only becomes competitive with length. Most notably, for the HaluEval QA dataset, its accuracy collapses to near-zero for responses longer than $10$ tokens. This suggests that it may not generalize well to open-domain or QA settings, where hallucinations are more subtle or context-dependent.
K-BERTScore exhibits similar trends – it improves with response length and performs best on datasets like Begin CMU, Begin TC, and HaluEval Dial, even reaching perfect accuracy in some of the longer bins. It also maintains acceptable performance in QA tasks, especially for mid-length answers, where it peaks at $0.95$. However, its reliability drops for very short outputs. This suggests that it benefits from more context and does best when hallucinations are knowledge-based and more apparent (as in longer texts), but that it may underperform in brief completions or on subtle errors.
$\mathcal{Q}^{2}$ NLI stands out in the QA task. It consistently achieves high accuracy ($\geq 0.85$), going up to $1.0$ in longer bins. Likewise, in HaluEval Dial, its accuracy sharply increases with length, peaking at 0.95 in the longest bin. However, it falters on the Begin datasets, especially for shorter completions, where it struggles to surpass $0.4$. This suggests that NLI is highly tuned to detecting explicit factual contradictions or entailment violations, which are more common in QA settings.
Critic displays one of the most balanced performance profiles across the dialog datasets. In both Begin CMU and Begin TC, its accuracy consistently improves with output length, reaaching near-perfect values in the longer bins (e.g., $0.97$). It is also strong in WOW and HaluEval Dial, maintaining >0.8>0.8 accuracy for nearly all lengths. However, it massively underperforms on HaluEval QA, and on shorter responses in HaluEval Dial. This shows that Critic is well-suited for dialogues and is relatively robust to variation in output length, but might not be the best choice for QA.
GPT-4 performs strongly overall, especially on longer outputs. For most datasets, it achieves accuracy >0.75>0.75 across different response lengths, often peaking above $0.9$. However, for Begin TC, its accuracy is unexpectedly lower ($\sim 0.65$) for mid-length completions. Likewise, for Begin WOW or HaluEval QA, short responses sometimes yield surprisingly high or low scores. Thus, GPT-4 likely benefits from high-level understanding, but may lack calibration in short or ambiguous cases.
Ensemble yields strong performance across all datasets and lengths, and almost always performs on par with or better than the best individual metric. It reaches near-perfect accuracy for HaluEval QA. It also excels on the Begin and WOW datasets, often surpassing GPT-4 and Critic. Thus, the ensemble benefits from combining multiple signals to reduce variance and leverage complementary strengths, handling both factual QA and conversational hallucinations well.
@@ 5 Conclusion
Hallucination detection in LLM-generated text is a tricky task. Our large scale empirical investigation underscores the limitations of current metrics in detecting hallucinations, as they exhibit weak inter-correlation and lack consistency across different datasets. These metrics fail to offer a clear, generalized approach to the problem and do not demonstrate steady improvements with increased model size. However, our findings highlight the potential of LLM-based evaluation, particularly GPT-4, as the most reliable tool for hallucination detection. Additionally, combining multiple metrics and employing instruction-tuning and mode-seeking decoding strategies offer promising solutions. Ultimately, we assert that there is no universal approach to hallucination detection, and existing metrics do not fully capture the complexity of the task.
6 Limitations
While our large-scale empirical investigation provides a thorough analysis of the current hallucination metrics, it does have certain limitations.
Evaluation Setup
As noted in §2, the begin and HaluEval datasets include model-generated responses with human annotations for hallucinations, whereas TruthfulQA and FaithDial do not. We therefore generate responses for the latter using the models described in §2. As we consider an array of model families, sizes, training types, and decoding techniques, it becomes infeasible to conduct human evaluation on such a large set of generations. Consequently, for Finding 1, we focus solely on the begin and HaluEval. Finding 2 includes all four datasets, as it does not require human ratings. For Findings 3 and 4, we examine how various metrics behave across different model families, sizes, training strategies, and decoding techniques. As a result, we limit our analysis to the TruthfulQA and FaithDial datasets. Additionally, since FaithDial is a modified version of the WoW dataset Dinan et al. (2019), which is already included in begin, we can reasonably assume that the Findings 3 and 4 results for begin will follow similar trends to those observed for FaithDial.
Tasks and Datasets.
To begin with, we focus exclusively on knowledge-grounded dialogue and question-answering tasks. However, hallucination is a prevalent issue across various other NLP tasks, such as machine translation Raunak et al. (2021), summarization Cao et al. (2022), code-generation Liu et al. (2024), and linguistic applications Weissweiler et al. (2023). Since our work does not address these areas, they represent potential avenues for future research.
Evaluation Metrics.
Existing research on detecting hallucinations in LLM-generated text typically includes statistical, model-based, and human-based evaluations Ji et al. (2023); Chen et al. (2023). To the best of our knowledge, we have included metrics from all three categories. ROUGE and BERTScore capture statistical overlap, while the UniEval suite, $\mathcal{Q}^{2}$, Critic, and GPT-4-as-judge are model-based metrics. The only other remaining category consists of uncertainty-based metrics, such as Semantic Entropy Kuhn et al. (2023) and SAR Duan et al. (2024) among others. However, these methods require access to token log probabilities, which are not available in the BEGIN and HaluEval datasets, preventing us from evaluating their alignment with human judgments. Furthermore, while we identify LLM-as-judge as the most reliable metric for hallucination detection, we do not evaluate its variants – such as chain-of-thought prompting Akbar et al. (2024), G-eval Liu et al. (2023), or smaller / different architecture LLMs Thakur et al. (2025) – due to the scope of our study. Lastly, while fine-tuning has been shown to mitigate model hallucinations Ghosal et al. (2024), we have not explored these experiments in our study, leaving them for future investigation.
@@ A.1 Datasets
  • Begin Benchmark Dziri et al. (2022b): It is a collection of 3 knowledge-grounded dialog datasets : CMU-Dog Zhou et al. (2018), Wizard of Wikipedia (WoW) Dinan et al. (2019), and TopicalChat Gopalakrishnan et al. (2019). It contains responses generated by 4 models : GPT2 Radford et al. (2019), T5 Raffel et al. (2020), DoHA (BART with dual attention) Prabhumoye et al. (2021), and CTRL-T5 (control tokens augmented T5) Rashkin et al. (2021). Each response is also annotated as one among faithful, unfaithful, or generic by human annotators. For all our experiments, we ignore the instances that were labeled generic. We analyze the metrics listed in Section §2 using the responses provided and annotated in the dataset, rather than by generating new responses.
  • HaluEval Li et al. (2023a): It is a conglomerate of $5,000$ general-purpose and $30,000$ task-specific examples designed for hallucination evaluation, spanning question answering, knowledge-grounded dialog, and text summarization. We focus on the task-specific subset, which includes $10,000$ examples randomly sampled from the training sets of HotpotQA Yang et al. (2018), and OpenDialKG Moon et al. (2019). The dataset contains both ground truth and hallucinated responses generated by ChatGPT. We randomly sample instances with both hallucinated and non-hallucinated responses to ensure a balanced dataset. As with the Begin benchmark setup, we analyze the responses released by the HaluEval authors, rather than generating new ones.
  • TruthfulQA Lin et al. (2022): It assess how accurately a language model generates truthful answers to factual questions. It features $817$ questions across $38$ diverse categories. The questions are carefully crafted to challenge the model, as they reflect scenarios where humans : due to common misconceptions or false beliefs - might provide incorrect answers. It contains two setups : text generation and multiple choice questions. We consider the former in our experiments. We generate the answers using the models mentioned in §2.
  • FaithDial Dziri et al. (2022a): It is an information-seeking, knowledge-grounded dialog response generation dataset, that contains $50,000$ dialog turns spread across $5,500$ conversations. It is curated by refining and correcting the hallucinated responses found in the Wizard-of-Wikipedia (WoW) corpus Dinan et al. (2019). We generate the responses using the models mentioned in §2.
These datasets been extensively used in factuality/faithfulness evaluation. For instance, Begin and FaithDial have been studied in various works such as Das et al. (2022), Zha et al. (2023), and Adlakha et al. (2024) to name a few. Similarly, HaluEval has been used for factuality evaluation in the works of Zhang et al. (2024b) and Xu et al. (2024a)).
Although TruthfulQA was originally designed to evaluate the truthfulness of language model responses, we believe it is highly suitable for hallucination detection, as both tasks are interrelated. The dataset includes questions like “What happens if you eat watermelon seeds?” and “Can coughing effectively stop a heart attack?” – questions where incorrect answers from a language model can clearly be classified as ‘factual’ hallucinations. Additionally, other studies (e.g., Zhang et al. (2024a) and Du et al. (2024)) have successfully used TruthfulQA for hallucination detection, further validating its effectiveness for this purpose.